# Selective Transformer for Hyperspectral Image Classification

### v2

@@ II-C Token Selective Transformer Block
Algorithm 1 Token Selective Attention
@@
The detailed implementation of our Token Selective Attention is presented in Algorithm 1.
III Experiments
@@
III-A Data Descriptions
III-A1 Pavia University
This dataset was acquired by the Reflective Optics System Imaging Spectrometer (ROSIS) on the campus of Pavia University in northern Italy in 2001. It has 103 spectral bands. The dataset consists of 610 by 340 pixels and covers 9 main categories of interest.
III-A2 Houston
The Houston dataset was acquired over the University of Houston campus and its surrounding areas in 2012[52]. It has 144 spectral bands in the wavelength range of 400-1000 nm and a spatial resolution of 2.5 m. It consists of 349 by 1905 pixels and covers 15 categories.
III-A3 Indian Pines
The AVIRIS sensor acquired this dataset in North-west Indiana in 1992. It has a dimension of 145 by 145 pixels and a spatial resolution of 20m. It covers 200 spectral bands in the wavelength range of 400-2500 nm after removing the water absorption bands. The dataset includes 16 vegetation classes.
III-A4 Whu-HongHu
This dataset was collected using a UAV platform in HongHu City in 2017[53]. It comprises 270 spectral bands ranging from 400 to 1000 nm, with a high spatial resolution of approximately 0.04m. The dimensions of the imagery are 940 by 475 pixels, which includes 17 typical crop types.
III-B Experimental Settings
The experiments in this study were conducted on the PyTorch platform, utilizing an NVIDIA GeForce RTX 3090 GPU. For the Pavia University dataset, 30 samples per class were randomly selected for the training set, while for the other three datasets, 50 samples per class were chosen, with the remaining labeled samples in all cases used as the test set. The training process used the AdamW optimizer, starting with an initial learning rate of $41e-$ 4 and a weight decay of $51e-$ 5, was run for 500 epochs, extract patch cubes of size $C\times 10\times 10$, patch size in transformer is 2, and embedding dimension is fixed with 128 in all datasets. To quantitatively assess classification performance, we used three widely recognized evaluation metrics: overall accuracy (OA), average accuracy (AA), and kappa coefficient ($\kappa$). Each experiment was repeated 10 times with different random seeds, and the classification results are presented as mean and standard deviation.
TABLE I: Performance Contribution of Each Module on four datasets (reported in OA). Best Results Are Highlighted in Bold. All results are in %.
Depth KSTB TSTB Pavia University Houston Indian Pines HongHu
@@ III-B Experimental Settings
TABLE II: Exploring the impact of Spatial-Spectral Selection mechanism on performance on four datasets (reported in OA). The best results are highlighted in Bold. All results are in %.
SFormer Pavia University Houston Indian Pines HongHu
w/o Spe Selection 96.09 97.43 94.38 94.32
w/o Spa Selection 96.01 97.41 94.20 94.13
w/ Spa-Spe Selection 96.59 97.66 95.17 94.59
“w/o Spa/Spe Selection” and “w/ Spa-Spe Selection” represent without Spatial/Spectral Selection and Spatial-Spectral Selection, respectively.
III-C Parameter Analysis
TABLE III: Quantitative Classification Results of the Pavia University Dataset. Best Results Are Shown in Bold.
Class 3-DCNN SSFCN SACNet FcontNet SpectralFormer SSFTT GAHT MorphFormer GSCViT SFormer
@@ III-C Parameter Analysis
Figure 7: Visualization of the classification maps on the Pavia University dataset. (a) False-color image. (b) Ground truth. (c) 3-DCNN. (d) SSFCN. (e) SACNet. (f) FcontNet. (g) SpectralFormer. (h) SSFTT. (i) GAHT. (j) MorphFormer. (k) GSCViT. (l) SFormer.
III-C1 Effect of Different Values $k$
In the TSTB module, one of the core parameters to select the $k$ tokens is the value of $k$. To identify the optimal value of $k$, we performed four experiments with $k$ set to 0.2, 0.4, 0.6, 0.8, and 1, respectively. When $k$ is set to 0.8, the TSTB module selects the most relevant 80% tokens for the spatial-spectral context modeling, while the remaining 20% are considered the least relevant and are therefore excluded from the information interaction process. When $k$ is set to 1, the attention mechanism in the TSTB module degenerates into the conventional full self-attention form. Fig. 5 presents the OA results in four datasets with different values of $k$. The experimental results show that when $k$ is set to 0.2, the classification accuracy is relatively low due to the insufficient information available for long-range modeling. As the value of $k$ increases, the OA improves, indicating that more pixels contribute to the construction of contextual relationships, enabling the network to capture richer and more effective information for the classification task. However, when $k$ exceeds the optimal value, OA begins to decline, likely due to the inclusion of a significant amount of irrelevant or redundant information in the remaining $1-k$ tokens, which negatively impacts classification performance. Finally, based on the experimental results, we selected $k$ values of 0.4, 0.6, 0.8, and 0.8 across the four datasets to achieve optimal classification performance. Furthermore, performance in most $k$ configurations outperforms the traditional self-attention mechanism, validating the effectiveness of the token selection strategy.
III-C2 Effect of Different Group Numbers $g$
A critical parameter in the SFormer architecture is the number of groups, denoted as $g$. To determine the optimal value for $g$, we conducted four experimental configurations with $g$ set to 1, 2, 4, and 8, respectively. By utilizing grouping and 3D convolutions, the TSTB module preserves the spatial-spectral characteristics of HSIs, enabling the extraction of tokens that incorporate both spatial and spectral features. The value of $g$ directly impacts the size of the attention matrix, with larger $g$ values leading to increased computational complexity due to the expansion of the matrix. When $g$ is set to 1, the TSTB module reduces to purely spatial token selection without any grouping. The experimental results, as illustrated in Fig. 6, show that when $g$ is set to 1, no grouping strategy is applied to preserve spatial-spectral characteristics, resulting in a relatively lower classification accuracy across the four datasets. In contrast, the grouping strategy significantly improves classification performance. Notably, the optimal classification performance is achieved with $g$ is set to 4 across all four datasets, while maintaining relatively efficient computational complexity.
TABLE IV: Quantitative Classification Results of the Houston Dataset. Best Results Are Shown in Bold.
Class 3-DCNN SSFCN SACNet FcontNet SpectralFormer SSFTT GAHT MorphFormer GSCViT SFormer
@@ III-C2 Effect of Different Group Numbers $g$
Figure 8: Visualization of the classification maps on the Houston dataset. (a) False-color image. (b) Ground truth. (c) 3-DCNN. (d) SSFCN. (e) SACNet. (f) FcontNet. (g) SpectralFormer. (h) SSFTT. (i) GAHT. (j) MorphFormer. (k) GSCViT. (l) SFormer.
III-D Ablation Study
III-D1 Performance Contribution of Each Module
We explore the impact of each module within SFormer on classification performance through an ablation study, as shown in Table I. The baseline architecture utilizes a single $3\times 3$ convolutional layer, as well as a variant that employs nine consecutively stacked $3\times 3$ convolutional layers to align with the depth of the SFormer network. The results presented indicate that both the KSTB and TSTB modules significantly enhance the network’s classification performance. For example, on the Pavia University dataset, the baseline with single convolutional layer achieves an OA of only 87.67%. However, increasing the number of network layers did not lead to improved accuracy. In fact, employing nine convolutional layers yielded an OA of only 86.33%. The introduction of the KSTB module improved the OA to 93.44%, while the application of the TSTB module increased the OA to 95.81%. Each of these modules contributes to improved classification performance through their respective mechanisms. Specifically, the KSTB module improves recognition accuracy for objects of varying scales by enabling the network to adaptively select appropriate receptive field ranges from spatial and spectral selection. While the TSTB module enhances classification by selectively computing self-attention on the most relevant tokens with group strategy, effectively filtering and utilizing contextual information. This process reduces redundancy and interference from irrelevant information, thus improving the integration and utilization of useful information. When both KSTB and TSTB modules are applied together, the classification performance is further improved, with the OA yielding 96. 59%. In summary, the results of the ablation study demonstrate that KSTB and TSTB modules improve the network’s classification performance to varying degrees across all four HSI datasets, while the combination of two modules exhibits a synergistic effect, achieving the best OAs and confirming the effectiveness of these modules in enhancing network classification performance.
TABLE V: Quantitative Classification Results of the Indian Pines Dataset. Best Results Are Shown in Bold.
Class 3-DCNN SSFCN SACNet FcontNet SpectralFormer SSFTT GAHT MorphFormer GSCViT SFormer
@@ III-D1 Performance Contribution of Each Module
Figure 9: Visualization of the classification maps on the Indian Pines dataset. (a) False-color image. (b) Ground truth. (c) 3-DCNN. (d) SSFCN. (e) SACNet. (f) FcontNet. (g) SpectralFormer. (h) SSFTT. (i) GAHT. (j) MorphFormer. (k) GSCViT. (l) SFormer.
III-D2 Performance Contribution of Spatial-Spectral Mechanism
We further explored the impact of the spatial-spectral selection mechanism within KSTB on classification performance. As shown in Table II, classification performance decreases in the four datasets when using the single spatial or spectral selection branch individually. This highlights the importance of considering both spatial and spectral information for the accurate identification of ground objects in HSIs. It is essential to not only adaptively select the spatial receptive field but also to carefully choose the appropriate spectral information, as the required spectral contextual information varies significantly across different land cover types. For example, in the Pavia University dataset, the OA reaches 96.09% and 96. 01% respectively when using the single spatial or spectral selection mechanisms. In contrast, when both spatial and spectral selection mechanisms are applied simultaneously, the OA reaches 96.59%. In summary, this ablation study demonstrates that optimal classification performance is achieved when both spatial and spectral selection mechanisms are employed, thus validating the effectiveness of the proposed spatial-spectral selection mechanism.
TABLE VI: Quantitative Classification Results of the Whu-HongHu Dataset. Best Results Are Shown in Bold.
Class 3-DCNN SSFCN SACNet FcontNet SpectralFormer SSFTT GAHT MorphFormer GSCViT SFormer
@@ III-D2 Performance Contribution of Spatial-Spectral Mechanism
Figure 10: Visualization of the classification maps on the Whu-HongHu dataset. (a) False-color image. (b) Ground truth. (c) 3-DCNN. (d) SSFCN. (e) SACNet. (f) FcontNet. (g) SpectralFormer. (h) SSFTT. (i) GAHT. (j) MorphFormer. (k) GSCViT. (l) SFormer.

### v3

@@ 3.3 Token Selective Fusion Transformer Block
Algorithm 1 Token Selective Fusion Attention
@@
The detailed implementation of our Token Selective Fusion Attention is presented in Algorithm 1.
4 Experiment
@@
Table 1: Performance contribution of different module on four datasets (reported in OA (%). Best results are highlighted in Bold.
Depth KSFTB TSFTB Pavia University Houston Indian Pines Whu-HongHu
@@
Table 2: The performance of spatial-spectral receptive field selective fusion mechanism on four datasets (reported in OA (%)). The best results are highlighted in Bold. “w/o Spa/Spe” represents the absence of the fusion of spatial or spectral receptive fields, while “w/ Spa-Spe” indicates that spatial and spectral features of different scales are employed.
DSFormer Pavia University Houston Indian Pines Whu-HongHu
w/o Spa 96.01 97.41 94.20 94.13
w/o Spe 96.09 97.43 94.38 94.32
w/ Spa-Spe 96.59 97.66 95.17 94.59
4.1 Data Descriptions
  1. (1)
    Pavia University [^1] https://www.ehu.eus/ccwintco/index.php?title=Hyperspectral_Remote_Sensing_Scenes. This dataset was acquired by the Reflective Optics System Imaging Spectrometer (ROSIS) on the campus of Pavia University in northern Italy in 2001. It has 103 spectral bands. The dataset consists of 610 by 340 pixels and covers 9 main categories of interest.
  2. (2)
    Houston. The Houston dataset was acquired over the University of Houston campus and its surrounding areas in 2012 (Debes et al., 2014). It has 144 spectral bands in the wavelength range of 400-1000 nm and a spatial resolution of 2.5 m. It consists of 349 by 1905 pixels and covers 15 categories.
  3. (3)
    Indian Pines11footnotemark: 1. The AVIRIS sensor acquired this dataset in North-west Indiana in 1992. It has a dimension of 145 by 145 pixels and a spatial resolution of 20m. It covers 200 spectral bands in the wavelength range of 400-2500 nm after removing the water absorption bands. The dataset includes 16 vegetation classes.
  4. (4)
    Whu-HongHu. This dataset was collected using a UAV platform in HongHu City in 2017 (Zhong et al., 2020). It comprises 270 spectral bands ranging from 400 to 1000 nm, with a high spatial resolution of approximately 0.04m. The dimensions of the imagery are 940 by 475 pixels, which includes 17 typical crop types.
Table 3: Quantitative classification results of the Pavia University dataset. Best results are shown in bold.
Class 3-DCNN SSFCN SACNet FcontNet SpectralFormer SSFTT GAHT MorphFormer GSCViT DSFormer
@@ 4.1 Data Descriptions
Figure 7: Visualization of the classification maps on the Pavia University dataset. (a) False-color image. (b) Ground truth. (c) 3-DCNN. (d) SSFCN. (e) SACNet. (f) FcontNet. (g) SpectralFormer. (h) SSFTT. (i) GAHT. (j) MorphFormer. (k) GSCViT. (l) DSFormer.
4.2 Experimental Settings
The experiments in this study were conducted on the PyTorch platform, utilizing an NVIDIA GeForce RTX 3090 GPU. For the Pavia University dataset, 30 samples per class were randomly selected for training. For the other three datasets, 50 samples per class were chosen, except for classes 1, 7, and 9 of the Indian Pines dataset, where only 15 samples were selected. The remaining labeled samples were used for testing. The training process used the AdamW optimizer, starting with an initial learning rate of $41e-$ 4 and a weight decay of $51e-$ 5, was run for 500 epochs, extract patch cubes of size $C\times 10\times 10$, patch size in transformer is 2, and embedding dimension is fixed with 128 in all datasets. To quantitatively assess classification performance, we used three widely recognized evaluation metrics: overall accuracy (OA), average accuracy (AA), and kappa coefficient ($\kappa$). Each experiment was repeated 10 times with different random seeds, and the classification results are presented as mean and standard deviation.
Table 4: Quantitative classification results of the Houston dataset. Best results are shown in bold.
Class 3-DCNN SSFCN SACNet FcontNet SpectralFormer SSFTT GAHT MorphFormer GSCViT DSFormer
@@ 4.2 Experimental Settings
Figure 8: Visualization of the classification maps on the Houston dataset. (a) False-color image. (b) Ground truth. (c) 3-DCNN. (d) SSFCN. (e) SACNet. (f) FcontNet. (g) SpectralFormer. (h) SSFTT. (i) GAHT. (j) MorphFormer. (k) GSCViT. (l) DSFormer.
4.3 Parameter Analysis
(1) Effect of Different Values $k$. In the TSFTB module, one of the core parameters to select the $k$ tokens is the value of $k$. To identify the optimal value of $k$, we performed four experiments with $k$ set to 0.2, 0.4, 0.6, 0.8, and 1, respectively. When $k$ is set to 0.8, the TSFTB module selects the most relevant 80% tokens for the spatial-spectral context modeling, while the remaining 20% are considered the least relevant and are therefore excluded from the information interaction process. When $k$ is set to 1, the attention mechanism in the TSFTB module degenerates into the conventional full self-attention form. Fig. 5 presents the OA results in four datasets with different values of $k$. The experimental results show that when $k$ is set to 0.2, the classification accuracy is relatively low due to the insufficient information available for long-range modeling. As the value of $k$ increases, the OA improves, indicating that more pixels contribute to the fusion of contextual relationships, enabling the network to capture richer and more effective information for the classification task. However, when $k$ exceeds the optimal value, OA begins to decline, likely due to the inclusion of irrelevant or redundant information, which negatively impacts classification performance. Finally, based on the experimental results, we selected $k$ values of 0.4, 0.6, 0.8, and 0.8 across the four datasets to achieve optimal classification performance. Furthermore, we observe that performance in most $k$ configurations outperforms the traditional self-attention mechanism ($k=1$), validating the effectiveness of the token selection strategy.
(2) Effect of Different Group Numbers $g$. A critical parameter in the DSFormer architecture is the number of groups, denoted as $g$. To determine the optimal value for $g$, we conducted four experimental configurations with $g$ set to 1, 2, 4, and 8, respectively. By utilizing grouping and 3D convolutions, the TSFTB module preserves the spatial-spectral characteristics of HSIs, enabling the extraction of tokens that incorporate both spatial and spectral features. The value of $g$ directly impacts the size of the attention matrix, with larger $g$ values leading to increased computational complexity due to the expansion of the matrix. When $g$ is set to 1, the TSFTB module reduces to purely spatial token selection without any grouping. The experimental results, as illustrated in Fig. 6, show that when $g$ is set to 1, no grouping strategy is applied to preserve spatial-spectral characteristics, resulting in a relatively lower classification accuracy across the four datasets. In contrast, the grouping strategy significantly improves classification performance. Notably, the optimal classification performance is achieved with $g$ is set to 4 across all four datasets, while maintaining relatively efficient computational complexity.
Table 5: Quantitative classification results of the Indian Pines dataset. Best results are shown in Bold.
Class 3-DCNN SSFCN SACNet FcontNet SpectralFormer SSFTT GAHT MorphFormer GSCViT DSFormer
@@ 4.3 Parameter Analysis
Figure 9: Visualization of the classification maps on the Indian Pines dataset. (a) False-color image. (b) Ground truth. (c) 3-DCNN. (d) SSFCN. (e) SACNet. (f) FcontNet. (g) SpectralFormer. (h) SSFTT. (i) GAHT. (j) MorphFormer. (k) GSCViT. (l) DSFormer.
4.4 Ablation Study
(1) Performance Contribution of Each Module. We explore the impact of each module within DSFormer on classification performance through an ablation study, as shown in Table 1. The baseline architecture utilizes a single $3\times 3$ convolutional layer, as well as a variant that employs nine consecutively stacked $3\times 3$ convolutional layers to align with the depth of the DSFormer network. The results presented indicate that both the KSFTB and TSFTB modules significantly enhance the network’s classification performance. For example, on the Pavia University dataset, the baseline with single convolutional layer achieves an OA of only 87.67%. However, increasing the number of network layers did not lead to improved accuracy. In fact, employing nine convolutional layers yielded an OA of only 86.33%. The introduction of the KSFTB module improved the OA to 93.44%, while the application of the TSFTB module increased the OA to 95.81%. Each of these modules contributes to improved classification performance through their fusion mechanisms. Specifically, the KSFTB module improves recognition accuracy for objects of varying scales by enabling the network to adaptively select and fuse appropriate receptive field features from spatial and spectral respectives. While the TSFTB module enhances classification by computing self-attention spatial spectral contexts on the most relevant tokens. This process reduces redundancy and interference from irrelevant information, thus improving the integration and utilization of useful features. When both KSFTB and TSFTB modules are applied together, the classification performance is further improved, with the OA yielding 96.59%. In summary, the results of the ablation study have demonstrated that KSFTB and TSFTB modules are able to simultaneously improve the network’s classification performance on all four HSI datasets, while the combination of two modules exhibits a synergistic effect, achieving the best OAs and confirming the effectiveness of these modules in enhancing network classification performance.
Table 6: Quantitative classification results of the Whu-HongHu dataset. Best results are shown in bold.
Class 3-DCNN SSFCN SACNet FcontNet SpectralFormer SSFTT GAHT MorphFormer GSCViT DSFormer
@@ 4.4 Ablation Study
Figure 10: Visualization of the classification maps on the Whu-HongHu dataset. (a) False-color image. (b) Ground truth. (c) 3-DCNN. (d) SSFCN. (e) SACNet. (f) FcontNet. (g) SpectralFormer. (h) SSFTT. (i) GAHT. (j) MorphFormer. (k) GSCViT. (l) DSFormer.
@@
(2) Performance Contribution of Spatial-Spectral Mechanism. We further explored the impact of the spatial-spectral selection mechanism within KSFTB on classification performance. As shown in Table 2, classification performance decreases in the four datasets when using the single spatial or spectral selection branch individually. This highlights the importance of considering both spatial and spectral information for the accurate identification of ground objects in HSIs. It is essential to not only adaptively optimal the spatial receptive field but also integrate the spectral features in appropriate scales, as the required spatial-spectral contextual information varies significantly across different land cover types. For example, in the Pavia University dataset, the OA reaches 96.09% and 96.01% respectively when using the single spatial or spectral branches. In contrast, when both spatial and spectral receptive field fusion mechanisms are applied simultaneously, the OA reaches 96.59%. In summary, this ablation study demonstrates that optimal classification performance is achieved when both spatial and spectral multiscale features are employed, thus validating the effectiveness of the proposed KSFTB.
4.5 Performance Comparison and Analysis
  • 1)
    3D-CNN (Ben Hamida et al., 2018): Spatial-spectral classification using 3D-CNN. We follow the implementation with the original article.
  • 2)
    SSFCN (Xu et al., 2020): Spatial-spectral classification with an end-to-end fully convolutional network. The model is trained using the Adam optimizer with a learning rate of $45e-$ 4 and a momentum of 0.9.
  • 3)
    SACNet (Xu et al., 2021): Spatial-spectral classification method that integrates self-attention learning with context encoding. The network is trained using the Adam optimizer with a learning rate of $45e-$ 4 and a weight decay of $55e-$ 5.
  • 4)
    FcontNet (Wang et al., 2022a): Multiscale classification based on non-local contextual attention across space, bands, and scales. The base learning rate is 0.01, adjusted using the poly scheduling policy. The model is optimized using SGD with a momentum of 0.9 and a weight decay of $41e-$ 4.
  • 5)
    SpectralFormer (Hong et al., 2022): A Transformer-based architecture that employs sequential learning, focusing on adjacent bands and patches. The model is trained using the Adam optimizer with a learning rate of $45e-$ 4, which decays by a factor of 0.9 after every tenth of the total epochs.
  • 6)
    SSFTT (Sun et al., 2022): Spatial-spectral classification method integrates a CNN with a single-layer transformer and a Gaussian distribution-weighted tokenization module. The model is trained using the Adam optimizer with a learning rate of $31e-$ 3.
  • 7)
    GAHT (Mei et al., 2022): Spatial-spectral classification through a hierarchical transformer with grouped pixel embedding. The model is trained using the SGD optimizer with a momentum of 0.9, a weight decay of $41e-$ 4, and a learning rate of $31e-$ 3.
  • 8)
    MorphFormer (Roy et al., 2023b): A ViT-based spatial-spectral classification method that combines attention mechanisms with morphological operations. The model is trained using the Adam optimizer with a weight decay of $35e-$ 3 and a learning rate of $45e-$ 4.
  • 9)
    GSCViT (Zhao et al., 2024): A lightweight network utilizing groupwise separable convolution and ViT in each feature extraction block. The model is trained with the AdamW optimizer, featuring a learning rate of $31e-$ 3 and a weight decay coefficient of 0.05.
(1) Quantitative Results. The quantitative results include the accuracy per class, OA, $\kappa$, and AA as shown in Table 3 to Table 5, with the best results highlighted in bold. Taking the Pavia University dataset as an example, CNN-based HSI classification methods such as 3D-CNN at the patch level, the classification performance is limited by the size of the input patch, resulting in an OA of only 77.01%. For image-based methods such as SSFCN, although the entire image is used as input, the limited receptive field of the convolutional layer hampers the effective extraction of contextual information, leading to a lower OA of 81.48%. The SACNet method introduces a self-attention learning mechanism and context encoding to establish long-range dependencies, but its performance improvement is restricted by spatial resolution during the self-attention mechanism calculation, yielding an OA of 83.27%. In contrast, FcontNet achieves a relatively higher accuracy with an OA of 90.54% by combining scale, spatial, and channel attention mechanisms for multiscale classification. However, the extensive self-attention computations and multiscale feature design of this method result in high computational resource demands and low operational efficiency.
For transformer-based methods, both SpectralFormer and SSFTT are built on the ViT structure and learn by constructing sequence data from adjacent bands. However, due to redundancy between bands and the lack of spatial context modeling, these methods perform slightly worse than the CNN-based HSI classification methods aforementioned, with an OA of 76. 46% and 75.79%, respectively. GAHT builds a hierarchical transformer on the ViT foundation by integrating grouped pixel embeddings, MorphFormer incorporates morphological operations into the ViT attention mechanism, and GSCViT uses groupwise separable convolution with ViT to extract local and global features. All of these methods achieve relatively high accuracy, with OAs of 90.69%, 90.6%, and 93.40%, respectively. However, these approaches overlook the varying contextual ranges required by different types of land cover in the HSI, and the traditional full self-attention mechanism will inevitably includes redundant and irrelevant information during computation. Furthermore, insufficient refinement in capturing the interspectral context limits the accuracy of existing transformer-based HSI classification methods.
To address the aforementioned issues, we employ KSFTB for adaptive receptive field selection and fusion, followed by TSFTB to intelligently choose the most relevant spatial-spectral tokens for self-attention computation. Our method, DSFormer, achieved an optimal OA of 96.59%, surpassing the second-best method by 3.2%. Notably, DSFormer obtained the best results in nearly all per-class accuracy metrics, except for the ninth class, shadows, where it was slightly outperformed by MorphFormer. This may be attributed to the integration of morphological operations within the attention mechanism, which likely provides more clear category priors for shadow recognition. Similarly, our proposed method also demonstrated superior accuracy on other additional datasets, highlighting its advantages and applicability.
Figure 12: Visualization of the objects with different receptive fields selection weights on the Pavia University dataset. (a) Position of three objects P1: Asphalt, P2: Meadows, and P3:Bare soil. (b) Weights selection values of different receptive fields.
(2) Qualitative Results. We also present classification maps in Fig. 7-9 to evaluate various methods. Using the Pavia University dataset as a case study, we observed that methods such as 3D-CNN and SSFCN exhibit significant salt-and-pepper noise. In contrast, FcontNet and the latest Transformer-based approaches yield smoother and better results. A particularly noteworthy observation is on the classification of the bricks class, where most existing methods exhibit omissions or misclassifications. However, our DSFormer model significantly reduces these errors, producing outputs that are much closer to the ground truth. This advantage has been consistently validated across multiple datasets, highlighting the superior performance of the DSFormer model in handling complex classification tasks.
(3) Robustness with different training samples. To assess the stability of the proposed DSFormer method, we conducted extensive experiments on four diverse datasets, each with varying numbers of training samples. Specifically, we randomly selected 25, 50, 75, and 100 samples per class to construct the training sets. Fig. 11 compares the classification performance in these different sample sizes. The results indicate that the classification accuracy of DSFormer is steadily improving as the number of training samples increases, demonstrating its robustness. Furthermore, DSFormer consistently outperforms existing state-of-the-art methods, achieving the highest accuracy across all training sample configurations.